# Phase-space localization at the lowest Landau level

### v1

@@ Abstract
We have thereby confirmed that a selection of states with strongly localized phase-space densities emerges when degenerate energy levels of free nonrelativistic bosons in a harmonic trap split due to weak contact interactions. The states with localized densities either lie at the bottom of the fine structure emanating from the unperturbed degenerate levels, or are obtained by acting on the lowest states at a different free-boson level with the center-of-mass boost $Z$. While, classically, being close to the energy minimum guarantees being severely restricted in phase space, quantum-mechanically, the presence of a potential well does not guarantee even the existence of bound states, let alone their strong localization. Our analysis thus demonstrates that the energy valleys defined by the saturation of the bound (18) are deep enough to provide for strong phase-space localization of the corresponding quantum energy eigenstates. It is an extra appealing feature that the structure of the LLL Hamiltonian can be exploited to give analytic expressions (12) for the wavefunctions of these special states [46]. By contrast, if we move to levels with $M>N$, where the lowest eigenstates of the fine structure are not ladder states, the degree of localization decreases, which correlates intuitively with the energy gap (binding energy) at the bottom of the spectrum becoming more shallow. We provide the corresponding plots in the appendix.
@@
An unusual aspect of our treatment is that we have used the phase space of the classical Schrödinger field, corresponding to second-quantized quantum mechanics, rather than the particle phase space (positions and momenta) of individual bosons. This choice played an essential role in our treatment. Indeed, had we chosen to work in the first-quantized representation, we would have had to deal with the phase space of, say, 70 particles in 2 dimensions, with all directions playing a significant role, pushing the dimensionality beyond what one could realistically work with [47]. By contrast, in the second quantized representation, the higher modes $\alpha_{n\gg 1}$ have very low occupation numbers and can be ignored in the analysis of phase-space densities, leaving behind a problem of moderately high, manageable dimension.
@@
We are not aware of prior treatments of the physics of ultracold atomic gases in the language of phase-space densities of the second-quantized representation. This way of handling quantum states is very common, on the other hand, in quantum optics [48], also in conjunction with experiments. We have demonstrated the utility of this picture for visualizing the properties of energy eigenstates of weakly interacting bosons. We expect this approach to have further fruitful applications.
Acknowledgements.
@@
Acknowledgments: This work has been supported by Research Foundation Flanders (FWO) through project G012222N, and by Vrije Universiteit Brussel through the Strategic Research Program High-Energy Physics. The work of MDC is partially supported by STFC consolidated grants ST/T000694/1 and ST/X000664/1 and by the Simons Investigator award ##\##620869. OE is supported by Thailand NSRF via PMU-B (grant number B13F670063). MP is supported by the FWO doctoral fellowship 1178725N. The computational resources and services used in this work were provided by the VSC (Flemish Supercomputer Center), funded by the Research Foundation Flanders (FWO) and the Flemish Government. In the numerics, we have used the partition generator [49].
Appendix
A derivation of the LLL Hamiltonian
@@
The conservation laws for the LLL Hamiltonian are
@@
Finally, we mention an additional structure controlling the energy eigenstates. It follows from the commutation relations (39) that the $Z$-operator acts as a raising operator on energy eigenstates, moving states from block $(N,M)$ to block $(N,M+1)$ and copying the $p_{N}(M)$ eigenvalues of $\mathcal{H}_{LLL}$ from the lower to the higher block. The LLL Hamiltonian can be diagonalized simultaneously with $ZZ^{\dagger}$, and the eigenvalues of this latter operator keep track of the number of times the raising operator $Z$ has been applied to the eigenstate. In an $(N,M)$-block, the eigenvalues of $ZZ^{\dagger}$ can straightforwardly be deduced from the commutation relations (39) and take values $-lN$, with $l\in\{0,1,\cdots,M\}$. The eigenvalues that are not coming from lower blocks are associated with null vectors of $ZZ^{\dagger}$.
Symmetries of the Husimi density
In the main text, we stated that the Husimi density of an eigenstate of an operator bilinear in the creation-annihilation operators is invariant under the symmetry transformations generated by the corresponding classical conserved quantity. We demonstrate the validity of this claim for infinitesimal transformations. Start by considering the quadratic operator
@@
To show (41), we first note that, to linear order in $\varepsilon$, $\hat{h}$ can be inserted at no cost in the Husimi density
@@
where we used that $\hat{h}$ annihilates the vacuum in going from line 1 to 2, and used the BCH formula a second time to group the exponentials in going from line 3 to 4. This proves (41). In the main text, our Hamiltonian commutes with both $N$ and $M$ so that the Husimi density of an energy eigenstate takes a constant value over each two-dimensional torus defined by these bilinears.
Details of the sampling procedure
@@
Once a maximum of the Husimi distribution is found, we run the Metropolis-Hastings algorithm with $\alpha_{\max}$ as the seed and a Gaussian target distribution constraining the classical values $H$, $N$, $M$ and $|Z|^{2}$ of the sampled points around their values at $\alpha_{\max}$. Our choice of the target distribution allows the Metropolis-Hastings algorithm to wander in phase space while remaining close to the shell of constant conservation laws $M_{\psi}$. The classical energy of the phase-space point where the Husimi distribution peaks is generally expected to be within quantum uncertainty of the quantum eigenstate energy. (Some differences between classical and quantum quantities may be attributed to ordering ambiguities.) The classical values of the other three conserved quantities, besides the Hamiltonian, turn out to be exactly equal to the corresponding quantum number. Regarding the standard deviations of the Gaussian distribution, one needs to compromise between a deviation large enough allowing the algorithm to efficiently move away from the two-dimensional torus $\Theta(\alpha_{\max})$ during the sampling, but also small enough to eliminate contamination of the samples with points where the Husimi density takes negligible values that would not provide any useful information on the topography of the Husimi density distribution. Since the Hamiltonian and $ZZ^{\dagger}$ operator are quartic in the modes, we let the standard deviations $\sigma_{H}$ and $\sigma_{|Z|^{2}}$ of the Gaussian distribution scale with $N$, while $\sigma_{N}$ and $\sigma_{M}$ are fixed at order 1, corresponding to the magnitude of typical ordering ambiguities in the corresponding operators. For definiteness, we chose the values $5\sqrt{2}\times(\sigma_{N},\sigma_{M},\sigma_{|Z|^{2}},\sigma_{H})=(1,1,N,N)$. In the Metropolis-Hastings sampling, the $(n\!+\!1)^{\text{th}}$ candidate sample point is generated by a (multivariate) Gaussian distribution, with the $n^{\text{th}}$ sample point taken as its mean. The standard deviation of this Gaussian is a measure for the step size to the next candidate, and was chosen to be
@@
Figure 2: Husimi function values over a cloud of phase-space points against the distance $d$ to the initial two-dimensional surface $\Theta(\alpha_{\max})$. Top: for the ladder state with $|Z|=0$, $N=30$ and $M=30$. Middle: for a non-ladder state with $|Z|=0$, $N=30$ and $M=30$. Bottom: for the non-ladder ground state of the Hilbert space block with $N=20$ and $M=30$. The Gaussian distribution used in the sampling algorithm has mean values determined by the classical values of the initial optimized phase-space point $(N_{\max},M_{\max},Z_{\max},H_{\max})=(30,30,0,225)$ (top), $(N_{\max},M_{\max},Z_{\max},H_{\max})=(30,30,0,240.32)$ (middle) and $(N_{\max},M_{\max},Z_{\max},H_{\max})=(20,30,0,84.60)$ (bottom). The standard deviations of the Gaussian distribution are set to $5\sqrt{2}\times(\sigma_{N},\sigma_{M},\sigma_{|Z|^{2}},\sigma_{H})=(1,1,N,N)$ for all three samplings. The three plots show again that the Husimi density is only well-localized on $\Theta(\alpha_{\max})$ when the considered eigenstate is a ladder state. Such localization is not observed when the ground state is not a ladder state.
$10^{-3}$. For numerical efficiency, we only vary the lowest 10 classical modes when moving from one sample to the next. This truncation does not significantly affect our results, as long as the modes $a_{n\geq 10}$ have small expectation values in the quantum state of interest.
No localization for non-ladder ground states
At $M>N$, ground states of the LLL system are not necessarily ladder states anymore. It is therefore interesting to compare the Husimi densities of ladder and non-ladder ground states. As an illustration, we perform the same analysis as done in the main text on
@@
(All of these states are annihilated by $Z^{\dagger}$.) The dimension of the $(30,30)$-block is $p_{N}(M)=5604$, while the $(20,30)$-block has dimension $p_{N}(M)=5507$. For the ($30,30$)-block, the ground state is still a ladder state and the top two plots of Fig. 2 show that the same differences between the ground state and a generic state can be observed as in the main text. For the ground state, the only phase-space points with large Husimi density are located in the vicinity of the two-dimensional surface $\Theta(\alpha_{\max})$. By contrast, for the generic eigenstate, we observe large Husimi values at phase-space points far away from $\Theta(\alpha_{\max})$.
For the ($20,30$)-block, the ground state is no longer a ladder state. In this case, no localization of the Husimi density is observed, as seen on the bottom plot of Fig. 2, where a swarm of points appears on the right of the main cloud, showing that there are points in phase space far away from the initial surface $\Theta(\alpha_{\max})$ where the Husimi function is nevertheless large. Thus, the non-ladder state is less tightly bound and less well-localized, compared to ladder states, despite being at the bottom of the spectrum.

### v2

@@ VI Conclusions
We have confirmed that a selection of states with strongly localized phase-space distributions emerges when degenerate energy levels of free nonrelativistic bosons in a harmonic trap split due to weak contact interactions. The states with localized distributions either lie at the bottom of the fine structure emanating from the unperturbed degenerate levels, or are obtained by acting on the lowest states at a different free-boson level with the center-of-mass boost $Z$. While, classically, being close to the energy minimum guarantees being severely restricted in phase space, quantum-mechanically, the presence of a potential well does not guarantee even the existence of bound states, let alone their strong localization. Our analysis thus demonstrates that the energy valleys defined by the saturation of the bound (19) are deep enough to provide for strong phase-space localization of the corresponding quantum energy eigenstates. It is an extra appealing feature that the structure of the LLL Hamiltonian can be exploited to give analytic expressions (13) for the wavefunctions of these special states [43]. By contrast, if we move to levels with $M>N$, where the lowest energy eigenstates of the fine structure are not ladder states, the degree of localization decreases, which correlates intuitively with the energy gap (binding energy) at the bottom of the spectrum becoming more shallow. We provide the corresponding plots in Appendix D.
@@
An unusual aspect of our treatment is that we have relied on the phase space of the classical Schrödinger field, corresponding to second-quantized quantum mechanics, rather than the particle phase space (positions and momenta) of individual bosons. This choice played an essential role in our treatment. Indeed, had we chosen to work in the first-quantized representation, we would have had to deal with the phase space of, say, 70 particles in 2 dimensions, with all directions playing a significant role, pushing the dimensionality beyond what one could realistically work with [44]. By contrast, in the second quantized representation, the higher modes $\alpha_{n\gg 1}$ have very low occupation numbers and can be ignored in the analysis of phase-space distributions, leaving behind a problem of moderately high, manageable dimension.
@@
We are not aware of prior treatments of the physics of ultracold atomic gases in the language of phase-space distributions of the second-quantized representation. This way of handling quantum states is very common, on the other hand, in quantum optics, also in conjunction with experiments. Indeed, the Schrödinger field plays the same role for identical nonrelativistic bosons that the Maxwell field plays for photons. Quantum phase-space distributions in terms of the Maxwell field oscillator modes have been used ubiquitously in quantum optics, see [45] for a textbook exposition. We have demonstrated the utility of this picture for visualizing the properties of energy eigenstates of weakly interacting bosons. This approach should have further fruitful applications.
Acknowledgments:
@@
This work has been supported by Research Foundation Flanders (FWO) through project G012222N, and by Vrije Universiteit Brussel through the Strategic Research Program High-Energy Physics. The work of MDC is partially supported by STFC consolidated grants ST/T000694/1 and ST/X000664/1 and by the Simons Investigator award ##\##620869. OE is supported by Thailand NSRF via PMU-B (grant number B13F670063). MP is supported by the FWO doctoral fellowship 1178725N. The computational resources and services used in this work were provided by the VSC (Flemish Supercomputer Center), funded by the Research Foundation Flanders (FWO) and the Flemish Government. In the numerics, we have used the partition generator [46].
Appendix A Derivation of the LLL Hamiltonian
@@
The conservation laws for the LLL Hamiltonian and their nonvanishing commutators are
@@
Finally, we mention an additional structure controlling the energy eigenstates. It follows from the commutation relations (40) that the $Z$-operator acts as a raising operator on energy eigenstates, moving states from block $(N,M)$ to block $(N,M+1)$ and copying the $p_{N}(M)$ eigenvalues of $\mathcal{H}_{LLL}$ from the lower to the higher block. The LLL Hamiltonian can be diagonalized simultaneously with $ZZ^{\dagger}$, and the eigenvalues of this latter operator keep track of the number of times the raising operator $Z$ has been applied to the eigenstate. In an $(N,M)$-block, the eigenvalues of $ZZ^{\dagger}$ can straightforwardly be deduced from the commutation relations (40) and take values $lN$, with $l\in\{0,1,\cdots,M\}$. The eigenvalues that are not coming from lower blocks are associated with null vectors of $ZZ^{\dagger}$.
Appendix B Symmetries of the Husimi distribution
In the main text, we stated that the Husimi distribution of an eigenstate of an operator bilinear in the creation-annihilation operators is invariant under the symmetry transformations generated by the corresponding classical conserved quantity. We demonstrate the validity of this claim for infinitesimal transformations. Start by considering the quadratic operator
@@ Appendix B Symmetries of the Husimi distribution
To show (42), we first note that, to linear order in $\varepsilon$, $\hat{h}$ can be inserted at no cost in the Husimi distribution
@@
where we used that $\hat{h}$ annihilates the vacuum in going from line 1 to 2, and used the BCH formula a second time to group the exponentials in going from line 3 to 4. This proves (42). In the main text, our Hamiltonian commutes with both $N$ and $M$ so that the Husimi distribution of an energy eigenstate takes a constant value over each two-dimensional torus defined by these bilinears.
Appendix C Details of the sampling procedure
@@
Once a maximum of the Husimi distribution is found, we run the Metropolis-Hastings algorithm with $\alpha_{\max}$ as the seed and a Gaussian target distribution constraining the classical values $H$, $N$, $M$ and $|Z|^{2}$ of the sampled points around their values at $\alpha_{\max}$. Our choice of the target distribution allows the Metropolis-Hastings algorithm to wander in phase space while remaining close to the shell of constant conservation laws $M_{\psi}$. The classical energy of the phase-space point where the Husimi distribution peaks is generally expected to be within quantum
@@
Figure 2: Husimi function values over a cloud of phase-space points against the distance $d$ to the initial two-dimensional surface $\Theta(\alpha_{\max})$. Top: for the ladder state with $|Z|=0$, $N=30$ and $M=30$. Middle: for a non-ladder state with $|Z|=0$, $N=30$ and $M=30$. Bottom: for the non-ladder ground state of the Hilbert space block with $N=20$ and $M=30$. The Gaussian distribution used in the sampling algorithm has mean values determined by the classical values of the initial optimized phase-space point $(N_{\max},M_{\max},Z_{\max},H_{\max})=(30,30,0,225)$ (top), $(N_{\max},M_{\max},Z_{\max},H_{\max})=(30,30,0,240.32)$ (middle) and $(N_{\max},M_{\max},Z_{\max},H_{\max})=(20,30,0,84.60)$ (bottom). The standard deviations of the Gaussian distribution are set to $5\sqrt{2}\times(\sigma_{N},\sigma_{M},\sigma_{|Z|^{2}},\sigma_{H})=(1,1,N,N)$ for all three samplings. The three plots show again that the Husimi distribution is only well-localized on $\Theta(\alpha_{\max})$ when the considered eigenstate is a ladder state. Such localization is not observed when the ground state is not a ladder state.
uncertainty of the quantum eigenstate energy. (Some differences between classical and quantum quantities may be attributed to ordering ambiguities.) The classical values of the other three conserved quantities, besides the Hamiltonian, turn out to be exactly equal to the corresponding quantum number. Regarding the standard deviations of the Gaussian distribution, one needs to compromise between a deviation large enough allowing the algorithm to efficiently move away from the two-dimensional torus $\Theta(\alpha_{\max})$ during the sampling, but also small enough to eliminate contamination of the samples with points where the Husimi distribution takes negligible values that would not provide any useful information on the topography of the Husimi distribution. Since the Hamiltonian and $ZZ^{\dagger}$ operator are quartic in the modes, we let the standard deviations $\sigma_{H}$ and $\sigma_{|Z|^{2}}$ of the Gaussian distribution scale with $N$, while $\sigma_{N}$ and $\sigma_{M}$ are fixed at order 1, corresponding to the magnitude of typical ordering ambiguities in the corresponding operators. For definiteness, we chose the values $5\sqrt{2}\times(\sigma_{N},\sigma_{M},\sigma_{|Z|^{2}},\sigma_{H})=(1,1,N,N)$. In the Metropolis-Hastings sampling, the $(n\!+\!1)^{\text{th}}$ candidate sample point is generated by a (multivariate) Gaussian distribution, with the $n^{\text{th}}$ sample point taken as its mean. The standard deviation of this Gaussian is a measure for the step size to the next candidate, and was chosen to be $10^{-3}$.   For numerical efficiency,   we only vary the lowest 10 classical modes when moving from one sample to the next. This truncation does not significantly affect our results, as long as the modes $a_{n\geq 10}$ have small expectation values in the quantum state of interest.
Appendix D No localization for non-ladder ground states
At $M>N$, ground states of the LLL system are not necessarily ladder states anymore. It is therefore interesting to compare the Husimi distributions of ladder and non-ladder ground states. As an illustration, we perform the same analysis as done in the main text on
@@ Appendix D No localization for non-ladder ground states
(All of these states are annihilated by $Z^{\dagger}$.) The dimension of the $(30,30)$-block is $p_{N}(M)=5604$, while the $(20,30)$-block has dimension $p_{N}(M)=5507$. For the ($30,30$)-block, the ground state is still a ladder state and the top two plots of Fig. 2 show that the same differences between the ground state and a generic state can be observed as in the main text. For the ground state, the only phase-space points with large Husimi distribution are located in the vicinity of the two-dimensional surface $\Theta(\alpha_{\max})$. By contrast, for the generic eigenstate, we observe large Husimi values at phase-space points far away from $\Theta(\alpha_{\max})$.
For the ($20,30$)-block, the ground state is no longer a ladder state. In this case, no localization of the Husimi distribution is observed, as seen on the bottom plot of Fig. 2, where a swarm of points appears on the right of the main cloud, showing that there are points in phase space far away from the initial surface $\Theta(\alpha_{\max})$ where the Husimi function is nevertheless large. Thus, the non-ladder state is less tightly bound and less well-localized, compared to ladder states, despite being at the bottom of the spectrum.